Find a general solution to the differential equation using the method of variation of parameters. y +10y + 25y = 3 e -50 T
Answer #1

y + 104+254= 3e5t = f(t) Armilary equation is mtiom +25=0 (m +50 m=-5,25 Complementary fumeting is Gente Let; Particular ante